OpenStudy (anonymous):

find the derivative of f(x)= (x +1)^1/2

OpenStudy (anonymous):

1/2(x+1)^(-1/2)

OpenStudy (anonymous):

????

OpenStudy (anonymous):

(1/2)(x+1)^(-1/2)

OpenStudy (anonymous):

The trick is.... (exponent)*(function)^(exponent - 1). So the derivative of f(x)=2x^2 is (2)*(2x)^(2-1) which equals 4x

OpenStudy (anonymous):

Yes i know that way but we cant do it that way we have to use the formula [f(x+h) - f(x)]/h

OpenStudy (anonymous):

The definition of the derivative... if f(x) = (x+1)^(1/2), then f(x+h) = (x+h+1)^(1/2). Plug both of these into your definition of the derivative. [(x+h+1)^(1/2) - (x+1)^(1/2)] / h You have to get rid of the square root. Do this by rationalizing. It's difficult to write how to do that in this text box. So let me give you a link. http://tutorial.math.lamar.edu/Classes/CalcI/DefnOfDerivative.aspx A problem very similar to yours is shown in Example 3

OpenStudy (anonymous):

yea i had them plugged in i just didnt know where to go from there